A conducting wire has a 1.0 mm diameter, a 2.0 m length, and a 50 mΩ resistance, What is the resistivity of the material?
1
Expert's answer
2019-01-15T09:02:11-0500

Resistivity of the material is:

ρ=RS/l=(R·πr^2)/l=(50·〖10〗^(-3)·3.14·〖(0.5)〗^2)/2=0.019625 (Ohm·〖mm〗^2)/m (where r=D/2=1 mm/2=0.5 mm)